OpenStudy (anonymous):

find the exact value of the expression for the given value of θ cos(2θ) if θ=pi/3

Parth (parthkohli):

If you like degrees, then\[{\pi \over 3} = {60}^{\circ}\]

Parth (parthkohli):

\[2\theta \implies 2\times 60 \implies 120\]

Parth (parthkohli):

Can you find cos(120 degrees), or should I draw a triangle?

OpenStudy (anonymous):

-0.5

OpenStudy (anonymous):

i think i can find cos(120)

Parth (parthkohli):

Good!

Parth (parthkohli):

But in degrees, not radians.

OpenStudy (anonymous):

i got 0.8142

Parth (parthkohli):

Let me check it for myself!

Parth (parthkohli):

@kizzfan911 You found the following:\[\cos(120 \text{ radians})\]But you have to find the following:\[\cos(\text{120 degrees)}\]

OpenStudy (anonymous):

how do i find it in degrees instead of radians?

OpenStudy (anonymous):

-.5 is what i got now. is that right?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

aight. thank you again.

Parth (parthkohli):

You're welcome :)